A piston contains 580 moles of an ideal monatomic
Answer #1

2 Cy 12-4?? CP-20 ?9 ? K mult kmle o6X Io T. T, Pi vl -(?06A105) (13) = T2-101. 64 k. =(580) (12.4 ) (101.6 y?2962) 28.62 K . R 5) (8.3) work =0 3 T-700 k. =530 x 20.74 x (Ho G-PMX106? 101,6

11) Work Done = Qin - Qout = 2.11 * 106 J

12) Efficiency = work/ Qin = (2.11*106) / ( 7.71 * 106)

= 0.27 or 27 %